# Optimistic Estimation of Convergence in Markov Chains with the Average-Mixing Time

### v1

@@ 1   Introduction
In this paper, we propose to address some of the above listed concerns about the worst-case $t_{\mathsf{mix}}$, by considering the average-mixing time $t_{\mathsf{mix}}^{\sharp}$ which recently came under the spotlight in (Espuny Díaz et al., 2022; Münch and Salez, 2022). Specifically, we will confirm that $t_{\mathsf{mix}}^{\sharp}$ is a more optimistic measure of convergence, in the sense where some chains can average-mix arbitrarily faster than they mix, even on a two-point space. Furthermore, we initiate the program of estimating $t_{\mathsf{mix}}^{\sharp}$ from a single trajectory and we demonstrate that it is statistically less costly than the problem pertaining to $t_{\mathsf{mix}}$. In fact, contrary to the worst-case problem, our analysis carries to a large subset of countable-state Markov chains. One should rightfully expect $t_{\mathsf{mix}}^{\sharp}$ to be less informative than $t_{\mathsf{mix}}$ since it quantifies a weaker notion of mixing. However, we confirm that knowledge of $t_{\mathsf{mix}}^{\sharp}$ remains of practical value in the sense where it can be used in known powerful decoupling techniques (Yu, 1994) and appear as a natural parameter controlling the large deviation rate for additive functionals. Along the way, we explore the connections between the average-mixing time and $\beta$-mixing, argued by Vidyasagar (2013); Mohri and Rostamizadeh (2010) to be “just the right” assumption for the analysis of machine learning algorithms with weakly-dependent data.
@@ 1.1   Related work
Average variants of the mixing time have recently made their apparition in the literature. For instance, Berestycki et al. (2018) analyzed random walks on the giant component of the Erdös–Rényi random graph $\mathcal{G}(n,p)$ with $p=\Theta(1/n)$, and showed that the mixing time when starting randomly from any vertex is the order $\mathcal{O}(\log n)$ while it was known to be of the order $\mathcal{O}(\log^{2}n)$ for the mixing time from a worst starting point. Following up, Espuny Díaz et al. (2022) extended the approach on a larger collection of graphs with small bottlenecks, and show applicability of their approach for the Newman–Watts small-world model and supercritically percolated expander graphs. In (Lovász and Kannan, 1999), the authors introduce the notion of a conductance averaged over subsets of states with various sizes, and demonstrate that it can lead to tighter convergence bounds for counting and volume computation problems. Finally, the average-mixing time $t^{\sharp}_{\mathsf{mix}}$, considered in (Münch and Salez, 2022) and the object of our study, is a measure of convergence of a Markov chain defined with respect to the chain being started from its stationary distribution (refer to Section 2), instead of the worst-case state.
@@ 1.2.1 Highlight the significance of the average-mixing time
we briefly demonstrate how the average-mixing time (together with an a priori known order of convergence) can be used to quantitatively capture the convergence rate of a Markov chain in lieu of the worst-case mixing time. Specifically, we show (Lemma 2.3) that the average-mixing time $t^{\sharp}_{\mathsf{mix}}(\delta)$ naturally appears in the sample complexity for controlling the deviation probability from 00 of a centered function $f\colon\mathcal{X}\to[-1,1]$. Namely, for $\frac{1}{n}\sum_{t=1}^{n}f(X_{t})\leq\varepsilon$ to hold with probability at least $1-\delta$, we show that under sub-exponential $\beta$-mixing of order $b$, it is sufficient to have a trajectory of
@@ 1.2.2 Estimation of $\beta$-mixing coefficients
For a general time-discrete, time-homogeneous Markov chain over a countable state space, we establish a parametric rate of convergence for its individual $\beta$-coefficients from a single stationary trajectory of observations $X_{1},\dots,X_{n}$. Namely, for $s\in\mathbb{N}$, we construct an estimator $\widehat{\beta}(s)$ for $\beta(s)$, and we show (Theorem 3.1) that
@@ 1.2.3 Estimation of the average-mixing time
with probability at least $1-\delta$. Furthermore, for finite spaces, setting $\xi=1/4$ for simplicity, the parameter $p$ can be chosen so as to obtain an upper bound (Corollary 4.2) of
@@
where $C\in\mathbb{R}_{+}$ is a universal constant. One can thus demonstrably estimate with high probability the average-mixing time sub-linearly in the number of states, answering a question posed by Luca Zanetti and John Sylvester (Zanetti and John, 2023). This discovery stands in stark contrast with the worst-case mixing time estimation problem, or the relaxation time estimation problem, for which the sample complexity necessarily depends inversely on the minimum stationary probability (Hsu et al., 2019; Wolfer and Kontorovich, 2024).
@@ 1.3   Outline
Section 2 introduces the average-mixing time, illustrates its significance and analyzes some of its properties. In Section 2.2, under geometric ergodicity and time-reversibility assumptions, we establish an upper bound of $t^{\sharp}_{\mathsf{mix}}$ in terms of the relaxation time of the chain, which is non-vacuous over countable state spaces and under suitable conditions on the stationary distribution. In Section 2.3 we recall some important connection between the average-mixing time and $\beta$-mixing properties of the chain.
Section 3 is dedicated to the problem of inferring the $\beta$-mixing coefficients and average-mixing time from a single trajectory of observations. In Section 3.1, we estimate the $\beta$-mixing coefficients of the chain; first individually and then as a sequence, with respect to the uniform norm. In Section 3.3, we specialize our analysis to the uniformly ergodic setting, and in Section 3.4 we deduce an upper bound for the problem of estimating the average-mixing time from a single trajectory under uniform ergodicity.
@@ 1.4   Notation and setting
We let $\mathcal{X}$ be a countably infinite space of symbols equipped (for simplicity) with a total order. Real functions on $\mathcal{X}$ are written as infinite row vectors and for $x\in\mathcal{X}$, $e_{x}$ is the function defined by $x^{\prime}\in\mathcal{X}$, $e_{x}(x^{\prime})=\delta\left[x=x^{\prime}\right]$. For $p\in(0,\infty)$, and a function $v\in\mathbb{R}^{\mathcal{X}}$,
@@
We write $\mathcal{W}(\mathcal{X})$ for the set of all infinite stochastic matrices over $\mathcal{X}$. For a finite trajectory of length $n\in\mathbb{N}$, started from an initial distribution $\mu\in\mathcal{P}(\mathcal{X})$ and evolving according to $P\in\mathcal{W}(\mathcal{X})$, the probability of observing $(x_{1},x_{2},\dots,x_{n})\in\mathcal{X}^{n}$ can be factored as
@@ 2   The average-mixing time
where $\pi$ is the stationary distribution of $P$. The above should be understood as the total variance distance of the chain to stationarity, averaged with respect to $\pi$ over the starting state. The limit $\lim_{t\to\infty}d^{\sharp}(t)=0$ always exists [^1] This follows from a connection between average-mixing time and $\beta$-mixing discussed at Section 2.3, thus for $\xi\in(0,1)$, the average-mixing time of $P$, defined (Münch and Salez, 2022) as
@@ 2.1   Average-mixing versus worst-case mixing
Following their respective definitions, it holds that $d^{\sharp}(t)\leq d(t)$. Thus, for any fixed $\xi\in(0,1)$, average-mixing is always faster than mixing. The equality $t_{\mathsf{mix}}=t^{\sharp}_{\mathsf{mix}}$ is notably achieved for transitive chains (Münch and Salez, 2022, Remark 1), which in particular include random walks on groups [^2] A chain is called transitive when for any $x_{0},x_{0}^{\prime}\in\mathcal{X}$, there exists a bijection $\phi\colon\mathcal{X}\to\mathcal{X}$, with $\phi(x_{0})=x_{0}^{\prime}$ and for all $x,x^{\prime}\in\mathcal{X},P(\phi(x),\phi(x^{\prime}))=P(x,x^{\prime})$.. In general however, the mixing time can be dramatically larger than the average-mixing time, even when the state space is small, as formalized below.
@@ 2.2   Spectral methods under reversibility and geometric ergodicity
The half quasi-norm of the stationary distribution $\|\pi\|_{1/2}$, making its first appearance in our study, is a natural measure of complexity of the stationary distribution $\pi$, directly related to its Rényi entropy of order $1/2$. Note that $\|\pi\|_{1/2}$ could be infinite, in which case Lemma 2.1 becomes vacuous. However, Lemma 2.1 compares favorably with the bound of Münch and Salez (2022, Theorem 5, Corollary 2), which is inversely proportional to the minimum (non-zero) transition probability $P_{\star}\doteq\min_{x,x^{\prime}\in\mathcal{X}}P(x,x^{\prime})$. Additionally, Lemma 2.1 can be instructively compared with Levin et al. (2009, Theorem 12.4) for the mixing time $t_{\mathsf{mix}}(\xi)$,
@@ 2.3   Connection with $\beta$-mixing
$\beta(\mathcal{A},\mathcal{B})\doteq\sup\frac{1}{2}\sum_{i=1}^{I}\sum_{j=1}^{J}\left|\mathbb{P}(A_{i}\cap B_{i})-\mathbb{P}(A_{i})\mathbb{P}(B_{j})\right|,$
@@
which depends only on the stationary properties of the chain. When there exists $\beta_{1},b\in\mathbb{R}_{+}$ such that for any $s\in\mathbb{N}$, $\beta(s)\leq\beta_{1}/s^{b}$, we say that the process $\beta$-mixes algebraically or polynomially. When there exists $\beta_{0},\beta_{1},b\in\mathbb{R}_{+}$ such that for any $s\in\mathbb{N}$, $\beta(s)\leq\beta_{0}e^{-\beta_{1}s^{b}}$ we say that the chain $\beta$-mixes sub-exponentially fast when $b\neq 1$, and exponentially fast when $b=1$. In Section 6, we will introduce a class of Markov transition operators, whose flavor of mixing (algebraic, sub-exponential, exponential) can be fully controlled. Notably (refer to the proof of Theorem 3.2), when $P$ is uniformly ergodic with mixing time $t_{\mathsf{mix}}$, it holds that
@@
hence the process is exponentially $\beta$-mixing with $\beta_{0}=2$ and $\beta_{1}=\log(2)/t_{\mathsf{mix}}$. However, nothing a priori precludes the $\beta$-mixing rate from being superior to (4). The connection between the average-mixing time and $\beta$-mixing provides additional motivation for the problem of estimating the $\beta$-mixing coefficients, whose knowledge is known to lead to powerful decoupling techniques (Yu, 1994). The problem of inferring mixing coefficients from a single trajectory of observations has recently received attention from the community. In the most recent work (Khaleghi and Lugosi, 2021), consistent estimators for $\beta$- and $\alpha$-mixing coefficients were proposed. In Section 3, we will provide more explicit and finer finite-sample bounds for the estimation problem in the specialized Markovian setting. A regime of particular interest to us is when the chain $\beta$-mixes much faster than $\Theta(t_{\mathsf{mix}})$. Henceforth in this manuscript, when they coincide under stationarity, we may use the notations $\beta$ and $d^{\sharp}$ interchangeably.
@@ 3.3   Estimation of $\beta$-mixing coefficients under uniform ergodicity
Under uniform ergodicity, the mixing time is finite, and the $\beta$-mixing coefficients decays exponentially. In this setting, we show how to obtain a sample complexity upper bound for high-probability estimates on the invividual $\beta$-mixing coefficients.
@@ 3.4   Estimation of the average-mixing time
We move on to the estimation of the average-mixing time. Since the average distance to stationarity $d^{\sharp}=\beta$ does not generally enjoy sub-multiplicativity (it does not hold that $t^{\sharp}_{\mathsf{mix}}(\xi)(1/4)$ can be bounded by $t^{\sharp}_{\mathsf{mix}}$ and a polynomial of $\log\frac{1}{\xi}$), we formalize desired guarantees somewhat differently from Hsu et al. (2019); Wolfer and Kontorovich (2024). Indeed, for the relaxation time $t_{\mathsf{rel}}$, we were seeking an estimator $\widehat{t_{\mathsf{rel}}}$ to multiplicative error $\varepsilon$. Namely, the estimators in Hsu et al. (2019); Wolfer and Kontorovich (2024) guarantee that with high probability,
@@ 4.1   Ergodic properties
{tikzcd}
@@
where for any $v\in\mathbb{R}^{\mathcal{X}}$, we wrote
$\left\|\mu\right\|_{V}\doteq\sup_{x\in\mathcal{X}}V(x)v(x).$
@@ 4.3   Infinite graphs with controlled growth
is the set of directed edges. The growth of a graph (Trofimov, 1985; Campbell et al., 2022) is defined as a function $g\colon\mathbb{N}\to\mathbb{N}\cup\left\{\infty\right\}$, where for $s\in\mathbb{N}$,
@@
where $\rho\colon\mathcal{X}\times\mathcal{X}^{2}\to\mathbb{N}\cup\left\{\infty\right\}$ denotes the of a directed graph, which is the maximum eccentricity of any vertex in the graph [^6] With the convention that a disconnected graph has infinite diameter., $(\mathcal{X},\mathcal{E})|_{\mathcal{S}}$ is the sub-graph of $(\mathcal{X},\mathcal{E})$ induced by the vertices in $\mathcal{S}$, and the supremum is taken over all $\mathcal{S}$ such that $(\mathcal{X},\mathcal{E})|_{\mathcal{S}}$ has diameter at most $s$.

### v2

@@ 1   Introduction
In this paper, we propose to address some of the above listed concerns about the worst-case $t_{\mathsf{mix}}$, by considering the average-mixing time $t_{\mathsf{mix}}^{\sharp}$ which recently came under the spotlight in (Díaz et al., 2024; Münch and Salez, 2023). Specifically, we will confirm that $t_{\mathsf{mix}}^{\sharp}$ is a more optimistic measure of convergence, in the sense where some chains can average-mix arbitrarily faster than they mix, even on a two-point space. Furthermore, we initiate the program of estimating $t_{\mathsf{mix}}^{\sharp}$ from a single trajectory and we demonstrate that it is statistically less costly than the problem pertaining to $t_{\mathsf{mix}}$. In fact, contrary to the worst-case problem, our analysis carries to a large subset of countable-state Markov chains. One should rightfully expect $t_{\mathsf{mix}}^{\sharp}$ to be less informative than $t_{\mathsf{mix}}$ since it quantifies a weaker notion of mixing. However, we confirm that knowledge of $t_{\mathsf{mix}}^{\sharp}$ remains of practical value in the sense where it can be used in known powerful decoupling techniques (Yu, 1994) and appear as a natural parameter controlling the large deviation rate for additive functionals. Along the way, we explore the connections between the average-mixing time and $\beta$-mixing, argued by Vidyasagar (2013); Mohri and Rostamizadeh (2010) to be “just the right” assumption for the analysis of machine learning algorithms with weakly-dependent data.
@@ 1.1   Related work
Average variants of the mixing time have recently made their apparition in the literature. For instance, Berestycki et al. (2018) analyzed random walks on the giant component of the Erdös–Rényi random graph $\mathcal{G}(n,p)$ with $p=\Theta(1/n)$, and showed that the mixing time when starting randomly from any vertex is the order $\mathcal{O}(\log n)$ while it was known to be of the order $\mathcal{O}(\log^{2}n)$ for the mixing time from a worst starting point. Following up, Díaz et al. (2024) extended the approach on a larger collection of graphs with small bottlenecks, and show applicability of their approach for the Newman–Watts small-world model and supercritically percolated expander graphs. In (Lovász and Kannan, 1999), the authors introduce the notion of a conductance averaged over subsets of states with various sizes, and demonstrate that it can lead to tighter convergence bounds for counting and volume computation problems. Finally, the average-mixing time $t^{\sharp}_{\mathsf{mix}}$, considered in (Münch and Salez, 2023) and the object of our study, is a measure of convergence of a Markov chain defined with respect to the chain being started from its stationary distribution (refer to Section 2), instead of the worst-case state.
@@ 1.2.1 Highlight the significance of the average-mixing time
we briefly demonstrate how the average-mixing time, together with an a priori known order of convergence, can be used to quantitatively capture the convergence rate of a Markov chain in lieu of the worst-case mixing time. Specifically, we show (Lemma 2.3) that the average-mixing time $t^{\sharp}_{\mathsf{mix}}(\delta)$ naturally appears in the sample complexity for controlling the deviation probability from zero of a centered function $f\colon\mathcal{X}\to[-1,1]$. Namely, for $\frac{1}{n}\sum_{t=1}^{n}f(X_{t})\leq\varepsilon$ to hold with probability at least $1-\delta$, we show that under sub-exponential $\beta$-mixing of order $b$, it is sufficient to have a trajectory of
@@ 1.2.2 Estimation of $\beta$-mixing coefficients
For a general time-discrete, time-homogeneous Markov chain over a countable state space, we establish a parametric rate of convergence for its individual $\beta$-coefficients from a single stationary trajectory of observations $X_{1},\dots,X_{n}$. Namely, for $s\in\mathbb{N}$, we construct an estimator $\widehat{\beta}(s)$ for $\beta(s)$, and we show (refer to Theorem 3.1) that the absolute mean deviation can be controlled as follows,
@@ 1.2.3 Estimation of the average-mixing time
with probability at least $1-\delta$. Furthermore, for finite spaces, setting $\xi=1/4$ for simplicity, the parameter $p$ can be chosen so as to obtain an upper bound (refer to Corollary 4.2) of
@@
where $C\in\mathbb{R}_{+}$ is a universal constant. One can thus demonstrably estimate with high probability the average-mixing time sub-linearly in the number of states, answering a question posed by Luca Zanetti and John Sylvester. This discovery stands in stark contrast with the worst-case mixing time estimation problem, or the relaxation time estimation problem, for which the sample complexity necessarily depends inversely on the minimum stationary probability (Hsu et al., 2019; Wolfer and Kontorovich, 2024).
@@ 1.3   Outline
Section 2 introduces the average-mixing time, illustrates its significance and analyzes some of its properties. In Section 2.2, under geometric ergodicity and time-reversibility assumptions, we establish an upper bound of $t^{\sharp}_{\mathsf{mix}}$ in terms of the relaxation time of the chain, which remains non-vacuous over countable state spaces and under suitable conditions on the stationary distribution. In Section 2.3 we recall some important connection between the average-mixing time and $\beta$-mixing properties of the chain.
Section 3 is dedicated to the problem of inferring the $\beta$-mixing coefficients and average-mixing time from a single trajectory of observations. In Section 3.1, we estimate the $\beta$-mixing coefficients of the chain; first individually and then as a sequence, with respect to the uniform norm. From Section 3.3, we specialize our analysis to the uniformly ergodic setting, and in Section 3.4 we deduce an upper bound for the problem of estimating the average-mixing time from a single trajectory.
@@ 1.4   Notation and setting
We let $\mathcal{X}$ be a countably infinite space of symbols equipped (for simplicity) with a total order. Real functions on $\mathcal{X}$ are written as infinite row vectors and for $x\in\mathcal{X}$, $e_{x}$ is the function defined by $x^{\prime}\in\mathcal{X}$, $e_{x}(x^{\prime})=\delta\left[x=x^{\prime}\right]$, where $\delta[\cdot]$ is the predicate function taking values in $\left\{0,1\right\}$. For $p\in(0,\infty)$, and a function $v\in\mathbb{R}^{\mathcal{X}}$,
@@
We write $\mathcal{W}(\mathcal{X})$ for the set of all infinite stochastic matrices over $\mathcal{X}$. For a finite trajectory of length $n\in\mathbb{N}$, started with an initial distribution $\mu\in\mathcal{P}(\mathcal{X})$ and evolving according to $P\in\mathcal{W}(\mathcal{X})$, the probability of observing $(x_{1},x_{2},\dots,x_{n})\in\mathcal{X}^{n}$ can be factored as
@@ 2   The average-mixing time
where $\pi$ is the stationary distribution of $P$. The above should be understood as the total variance distance of the chain to stationarity, averaged with respect to $\pi$ over the starting state. The limit $\lim_{t\to\infty}d^{\sharp}(t)=0$ always exists [^1] This fact follows from a connection between average-mixing time and $\beta$-mixing discussed at Section 2.3., thus for $\xi\in(0,1)$, the average-mixing time of $P$, defined (Münch and Salez, 2023) as
@@ 2.1   Average-mixing versus worst-case mixing
Following their respective definitions, it holds that $d^{\sharp}(t)\leq d(t)$. Thus, for any fixed $\xi\in(0,1)$, average-mixing occurs always faster than mixing. The equality $t_{\mathsf{mix}}=t^{\sharp}_{\mathsf{mix}}$ is notably achieved for transitive chains (Münch and Salez, 2023, Remark 1), which in particular include random walks on groups [^2] A chain is called transitive when for any $x_{0},x_{0}^{\prime}\in\mathcal{X}$, there exists a bijection $\phi\colon\mathcal{X}\to\mathcal{X}$, with $\phi(x_{0})=x_{0}^{\prime}$ and for all $x,x^{\prime}\in\mathcal{X},P(\phi(x),\phi(x^{\prime}))=P(x,x^{\prime})$.. In general however, the mixing time can be dramatically larger than the average-mixing time, even when the state space is small, as formalized below.
@@ 2.2   Spectral methods under reversibility and geometric ergodicity
The half quasi-norm of the stationary distribution $\|\pi\|_{1/2}$, making its first appearance in our study, is a natural measure of complexity of the stationary distribution $\pi$, directly related to its Rényi entropy of order $1/2$. Note that $\|\pi\|_{1/2}$ could be infinite, in which case Lemma 2.1 becomes vacuous. However, Lemma 2.1 compares favorably with the bound of Münch and Salez (2023, Theorem 5, Corollary 2), which is inversely proportional to the minimum (non-zero) transition probability $P_{\star}\doteq\min_{x,x^{\prime}\in\mathcal{X}}P(x,x^{\prime})$. Additionally, Lemma 2.1 can be instructively compared with Levin et al. (2009, Theorem 12.4) for the mixing time $t_{\mathsf{mix}}(\xi)$,
@@ 2.3   Connection with $\beta$-mixing
$\beta(\mathcal{A},\mathcal{B})\doteq\sup\frac{1}{2}\sum_{i=1}^{I}\sum_{j=1}^{J}\left|\mathbb{P}(A_{i}\cap B_{j})-\mathbb{P}(A_{i})\mathbb{P}(B_{j})\right|,$
@@
which depends only on the stationary properties of the chain. When there exists $\beta_{1},b\in\mathbb{R}_{+}$ such that for any $s\in\mathbb{N}$, $\beta(s)\leq\beta_{1}/s^{b}$, we say that the process $\beta$-mixes algebraically or polynomially. When there exists $\beta_{0},\beta_{1},b\in\mathbb{R}_{+}$ such that for any $s\in\mathbb{N}$, $\beta(s)\leq\beta_{0}e^{-\beta_{1}s^{b}}$ we say that the chain $\beta$-mixes sub-exponentially when $b\neq 1$, and exponentially when $b=1$. In Section 6, we will introduce a class of Markov transition operators, whose flavor of mixing (algebraic, sub-exponential, exponential) can be fully controlled. Notably (refer to the proof of Theorem 3.2), when $P$ is uniformly ergodic with mixing time $t_{\mathsf{mix}}$, it holds that
@@
hence the process is exponentially $\beta$-mixing with $\beta_{0}=2$ and $\beta_{1}=\log(2)/t_{\mathsf{mix}}$. However, nothing a priori precludes the $\beta$-mixing rate from being superior to (4). The connection between the average-mixing time and $\beta$-mixing provides additional motivation for the problem of estimating the $\beta$-mixing coefficients, whose knowledge is known to lead to powerful decoupling techniques (Yu, 1994). The problem of inferring mixing coefficients from a single trajectory of observations has recently received attention from the community. In the most recent work (Khaleghi and Lugosi, 2023), consistent estimators for $\beta$- and $\alpha$-mixing coefficients were proposed. In Section 3, we will provide more explicit and finer finite-sample bounds for the estimation problem in the specialized Markovian setting. A regime of particular interest to us is when the chain $\beta$-mixes much faster than $\Theta(t_{\mathsf{mix}})$. Henceforth in this manuscript, when they coincide under stationarity, we may use the notations $\beta$ and $d^{\sharp}$ interchangeably.
@@ 3.3   Estimation of $\beta$-mixing coefficients under uniform ergodicity
Under uniform ergodicity, the mixing time is finite, and the $\beta$-mixing coefficients decays exponentially. In this setting, we show how to obtain a sample complexity upper bound for high-probability estimates on the individual $\beta$-mixing coefficients.
@@ 3.4   Estimation of the average-mixing time
We move on to the estimation of the average-mixing time. Since the average distance to stationarity $d^{\sharp}=\beta$ does not generally enjoy sub-multiplicativity (it does not hold that $t^{\sharp}_{\mathsf{mix}}(\xi)$ can be bounded by $t^{\sharp}_{\mathsf{mix}}(1/4)$ and a polynomial of $\log\frac{1}{\xi}$), we formalize desired guarantees somewhat differently from Hsu et al. (2019); Wolfer and Kontorovich (2024). Indeed, for the relaxation time $t_{\mathsf{rel}}$, we were seeking an estimator $\widehat{t_{\mathsf{rel}}}$ to multiplicative error $\varepsilon$. Namely, the estimators in Hsu et al. (2019); Wolfer and Kontorovich (2024) guarantee that with high probability,
@@ 4.1   Ergodic properties
Bound on ${\text{Bound on }\mathcal{J}_{p}^{(s)}\text{ uniform over }s}$ Condition 1Condition 2Condition 3Bound on ${\text{Bound on }\mathcal{J}_{p}^{(s)}\text{ sub-linear in }s}$Corollary 4.1Corollary 4.1Lemma 4.2Lemma 4.3Lemma 4.1
@@
where for any $\nu\in\mathbb{R}^{\mathcal{X}}$, we wrote
$\left\|\nu\right\|_{V}\doteq\sup_{x\in\mathcal{X}}V(x)\nu(x).$
@@ 4.3   Infinite graphs with controlled growth
is the set of directed edges. The growth of a graph (Trofimov, 1985; Campbell et al., 2023) is defined as a function $g\colon\mathbb{N}\to\mathbb{N}\cup\left\{\infty\right\}$, where for $s\in\mathbb{N}$,
@@
where $\rho\colon\mathcal{X}\times\mathcal{X}^{2}\to\mathbb{N}\cup\left\{\infty\right\}$ denotes the diameter of a directed graph, which is the maximum eccentricity of any vertex in the graph [^6] With the convention that a disconnected graph has infinite diameter., $(\mathcal{X},\mathcal{E})|_{\mathcal{S}}$ is the sub-graph of $(\mathcal{X},\mathcal{E})$ induced by the vertices in $\mathcal{S}$, and the supremum is taken over all $\mathcal{S}$ such that $(\mathcal{X},\mathcal{E})|_{\mathcal{S}}$ has diameter at most $s$.